A) (a) 4 B (a)3 C) (a) 3 (b) 9.5 (c)7.5-11.5 (b) 9.5 (c) 7.5-11.5 (b) 10.5 (c) 8-11, 5 D) (a) 4 (b) 10.5 (c) 8-11 ID: ES6L 2.1.1-2 Objective: (2.1) Interpret a Frequency Distribution rovide an appropriate response. 12) Use the ogive below to approximate the cumulative frequency for 24 hours. 12) Leisure Time of College Students E 70 60 50 a 40 30 20 S 10 9.5 14.5 19.5 245 29.5 345 39.5 Number of hours A)75 ID: ES6L 2.1.2-2 Objective: (2.1) Interpret Frequency Distribution Graphs B) 27 C) 63 D) 17 sample of candies have weights that vary from 2.35 grams to 4.75 grams. Use this information to 13)_D d the upper and lower limits of the first class if you wish to construct a frequency distribution with 12 classes. A) 2.35-2.65 B) 2.35-2.75 C) 2.35-2.54 2.35-2.55
Answer #1

12.) From the graph ,we can observe that the approximate cumulative frequency of :

19.5 is between 30 to 40

and

24.5 is between 65 to 70

so we can say that the cumulative fequency of 24 hours will be slightly less than 70.

By looking on the options 63 is the best fit

Ans: (C)

13.) Least weight limit of candies = 2.35 grams

Upper weight limit of candies = 4.75 grams

We want 12 classes in the frequency table

\small \therefore The width of each interval = (4.75-2.35)/12

= 0.2

So, the first interval will be = 2.35 - (2.35+0.2)

= 2.35 - 2.55

Ans: (D)